# Spec2SVA-Eval: Evaluating LLMs for Assertion Generation from Design Specification

### v3

@@ I Introduction
During the verification process, assertion-based verification (ABV) [1] is a widely adopted technique, which utilizes assertions crafted from specifications to verify the functional behavior of RTL designs. ABV can be conducted either through simulation with testbenches or using formal property verification (FPV) techniques. Temporal logic, particularly SystemVerilog Assertions (SVA), is commonly employed for specifying these properties. However, a significant challenge in ABV is the generation of sufficient, high-quality assertions. Currently, designing SVAs manually is a time-consuming and error-prone task, demanding unignorable human effort.
To address this challenge, research has focused on generating SVAs automatically. The automatic approaches can be categorized into two types: dynamic mining from simulation traces and static analysis of specifications. Dynamic methods [2, 3, 4] generate assertions by combining simulating test traces and static analysis of design constraints. However, a critical limitation of dynamic methods is that both the generation and evaluation of assertions are on the same RTL design without referring to a golden reference model. This could lead to the generation of incorrect SVAs due to flaws in the RTL design, which these methods might not detect. On the other hand, existing static methods depend either on the pre-defined templates [5, 6] or on machine learning (ML) technologies [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. The template-based methods also require a deep understanding of the design function to fill in the templates. As for the ML-based methods, they explore both traditional natural language processing (NLP) and emerging Generative AI techniques like Large Language Models (LLMs). We further categorize the existing static ML-based methods based on their application in different design phases: the RTL and pre-RTL stages.
Table I details these ML-based SVA generation methods in both the RTL stage and the pre-RTL stage. During the RTL stage, the process typically involves using LLMs to process both human-written specification sentences and the RTL design to generate SVAs describing security or functional properties [7, 8, 9]. However, similar to the dynamic methods, inaccuracies in RTL implementations could result in flawed SVAs.
@@
RTL [7] LLM-based Sentences from Engineers Security
[8, 9] Function
Pre-RTL [10, 11, 12, 13, 14, 15, 16, 17] NLP-based
@@
  • Work [8] only open-source the result on a FIFO, while work [9] only demonstrate based on tiny designs such as FSM and DFF.
@@
When it comes to the pre-RTL stage, with the natural language specification document finalized, RTL designers proceed to implement behavior satisfying this golden specification. Numerous studies [10, 11, 12, 13, 14, 15, 16, 17] have employed NLP techniques to generate SVAs from sentences extracted by humans. These works focused on processing sentences identified from a comprehensive document of specification. However, specification extraction requires tremendous human efforts, and the NLP-based generation process faces challenges in generalizing across diverse grammatical variations. Additionally, the evaluation of SVAs generated through these methods typically depends on design-specific checkers, such as protocol and processor checkers, and is therefore difficult to extend to other design types.
@@ II-B LLM for EDA
Recent advancements in LLMs like ChatGPT [18] have not only demonstrated remarkable capability in content generation but also evolved to assist humans in various roles as agents. The application of LLMs in the field of electronic design automation (EDA) is an emerging area of exploration. Besides employing LLMs for assertion generation [7, 8, 9], recent studies have employed LLMs for tasks such as RTL code generation [19, 20, 21, 22, 23, 24, 25, 26] and syntax correction [27]. Additionally, LLM-based solutions have been developed to facilitate the interaction with EDA tools [26, 28], design architecture for AI accelerators [29, 30], fix security bugs [31], generate and review specification documents [32], etc. These diverse applications and research efforts indicate a highly promising future for LLMs in enhancing and revolutionizing chip design processes.
@@ III-B Specification Information Extraction
Facing this challenge, existing methods can only deal with sentence-level specifications. Some works [14, 16] utilize ML methods to assess the relevance of manually extracted specification sentences to the intended assertions. Other studies [10, 11, 12, 15, 17] directly use the human-identified or human-written assertion-related sentences. This reliance on sentence-level analysis limits the ability of the above NLP-based methods to fully automate SVA generation for practical applications
@@ III-D Automatic Assertion Generation
Considering the precise syntax and writing rules inherent to SVAs and the potential for the original LLM failing to generate syntactically correct SVAs, as discussed in [8], we incorporate the Retrieval Augmented Generation (RAG) technique to enhance the LLM’s capability for SVA generation. This approach is enriched by a knowledge database comprising tutorials and textbooks on SVA and formal property verification [33, 34, 35], providing a robust foundation for the LLM to access and retrieve relevant SVA knowledge based on the input query, thereby enhancing the quality of the generated SVAs.
@@ III-E Generated Assertion Evaluation
After the SVAs are generated, evaluating their effectiveness is crucial. While some previous studies such as [12, 14] suggest using specific property checkers for this purpose, such an approach is limited to particular design types like protocols and processors and lacks generalizability to all VLSI designs. Other methods like [8] involve manual verification by engineers using EDA tools, which is vulnerable to human error.
@@ IV-D Ablation Study
For the SVA quality of commercial solutions, the original GPT-4 model tends to produce SVAs with syntax errors, similar to observations made in the previous study [8]. This issue is addressed in AssertLLM through the application of RAG techniques, which enrich the model with specific knowledge on SVA and Formal Property Verification (FPV). In the evaluation results, GPT-4 failed to generate any correct SVAs for IO ports and only succeeded in creating accurate reset check assertions for registers, leading to an overall correct proportion of only 8%.
@@ V-A Coverage in SVA Evaluation
Some works [8] propose to leverage the coverage metric, especially the cone-of-influence (COI) coverage to evaluate the quality of generated SVAs. COI coverage relies on analyzing the signals exercised during simulation or formal verification, which significantly involves internal signals within the design.

### v4

# AssertLLM: Generating and Evaluating Hardware Verification Assertions from Design Specifications via Multi-LLMs

@@ I Introduction
During the verification process, assertion-based verification (ABV) [witharana2022survey] is a widely adopted technique, which utilizes assertions crafted from specifications to verify the functional behavior of RTL designs. ABV can be conducted either through simulation with testbenches or using formal property verification (FPV) techniques. Temporal logic, particularly SystemVerilog Assertions (SVA), is commonly employed for specifying these properties. However, a significant challenge in ABV is the generation of sufficient, high-quality assertions. Currently, designing SVAs manually is a time-consuming and error-prone task, demanding unignorable human effort.
To address this challenge, research has focused on generating SVAs automatically. The automatic approaches can be categorized into two types: dynamic mining from simulation traces and static analysis of specifications. Dynamic methods [germiniani2022harm, danese2017team, vasudevan2010goldmine] generate assertions by combining simulating test traces and static analysis of design constraints. However, a critical limitation of dynamic methods is that both the generation and evaluation of assertions are on the same RTL design without referring to a golden reference model. This could lead to the generation of incorrect SVAs due to flaws in the RTL design, which these methods might not detect. On the other hand, existing static methods depend either on the pre-defined templates [orenes2021autosva, fang2023r] or on machine learning (ML) technologies [kande2023llm, orenes2023using, sun2023towards, harris2016glast, krishnamurthy2019controlled, zhao2019automatic, krishnamurthy2019ease, frederiksen2020automated, keszocze2019chatbot, parthasarathy2021spectosva, aditi2022hybrid]. The template-based methods also require a deep understanding of the design function to fill in the templates. As for the ML-based methods, they explore both traditional natural language processing (NLP) and emerging Generative AI techniques like Large Language Models (LLMs). We further categorize the existing static ML-based methods based on their application in different design phases: the RTL and pre-RTL stages.
Table I details these ML-based SVA generation methods in both the RTL stage and the pre-RTL stage. During the RTL stage, the process typically involves using LLMs to process both human-written specification sentences and the RTL design to generate SVAs describing security or functional properties [kande2023llm, orenes2023using, sun2023towards]. However, similar to the dynamic methods, inaccuracies in RTL implementations could result in flawed SVAs.
@@
RTL [kande2023llm] LLM-based Sentences from Engineers Security
[orenes2023using, sun2023towards] Function
Pre-RTL [harris2016glast, krishnamurthy2019controlled, zhao2019automatic, krishnamurthy2019ease, frederiksen2020automated, keszocze2019chatbot, parthasarathy2021spectosva, aditi2022hybrid] NLP-based
@@
  • Work [orenes2023using] only open-source the result on a FIFO, while work [sun2023towards] only demonstrate based on tiny designs such as FSM and DFF.
@@
When it comes to the pre-RTL stage, with the natural language specification document finalized, RTL designers proceed to implement behavior satisfying this golden specification. Numerous studies [harris2016glast, krishnamurthy2019controlled, zhao2019automatic, krishnamurthy2019ease, frederiksen2020automated, keszocze2019chatbot, parthasarathy2021spectosva, aditi2022hybrid] have employed NLP techniques to generate SVAs from sentences extracted by humans. These works focused on processing sentences identified from a comprehensive document of specification. However, specification extraction requires tremendous human efforts, and the NLP-based generation process faces challenges in generalizing across diverse grammatical variations. Additionally, the evaluation of SVAs generated through these methods typically depends on design-specific checkers, such as protocol and processor checkers, and is therefore difficult to extend to other design types.
@@ II-B LLM for EDA
Recent advancements in LLMs like ChatGPT [achiam2023gpt] have not only demonstrated remarkable capability in content generation but also evolved to assist humans in various roles as agents. The application of LLMs in the field of electronic design automation (EDA) is an emerging area of exploration. Besides employing LLMs for assertion generation [kande2023llm, orenes2023using, sun2023towards], recent studies have employed LLMs for tasks such as RTL code generation [liu2023rtlcoder, blocklove2023chip, lu2023rtllm, liu2023verilogeval, thakur2023benchmarking, thakur2023autochip, nair2023generating, liu2023chipnemo] and syntax correction [tsai2023rtlfixer]. Additionally, LLM-based solutions have been developed to facilitate the interaction with EDA tools [liu2023chipnemo, he2023chateda], design architecture for AI accelerators [fu2023gpt4aigchip, yan2023viability], fix security bugs [ahmad2023fixing], generate and review specification documents [li2024specllm], etc. These diverse applications and research efforts indicate a highly promising future for LLMs in enhancing and revolutionizing chip design processes.
@@ III-B Specification Information Extraction
Facing this challenge, existing methods can only deal with sentence-level specifications. Some works [frederiksen2020automated, parthasarathy2021spectosva] utilize ML methods to assess the relevance of manually extracted specification sentences to the intended assertions. Other studies [harris2016glast, krishnamurthy2019controlled, zhao2019automatic, keszocze2019chatbot, aditi2022hybrid] directly use the human-identified or human-written assertion-related sentences. This reliance on sentence-level analysis limits the ability of the above NLP-based methods to fully automate SVA generation for practical applications
@@ III-D Automatic Assertion Generation
Considering the precise syntax and writing rules inherent to SVAs and the potential for the original LLM failing to generate syntactically correct SVAs, as discussed in [orenes2023using], we incorporate the Retrieval Augmented Generation (RAG) technique to enhance the LLM’s capability for SVA generation. This approach is enriched by a knowledge database comprising tutorials and textbooks on SVA and formal property verification [seligman2023formal, mehta2020systemverilog, vijayaraghavan2005practical], providing a robust foundation for the LLM to access and retrieve relevant SVA knowledge based on the input query, thereby enhancing the quality of the generated SVAs.
@@ III-E Generated Assertion Evaluation
After the SVAs are generated, evaluating their effectiveness is crucial. While some previous studies such as [zhao2019automatic, frederiksen2020automated] suggest using specific property checkers for this purpose, such an approach is limited to particular design types like protocols and processors and lacks generalizability to all VLSI designs. Other methods like [orenes2023using] involve manual verification by engineers using EDA tools, which is vulnerable to human error.
@@ IV-D Ablation Study
For the SVA quality of commercial solutions, the original GPT-4 model tends to produce SVAs with syntax errors, similar to observations made in the previous study [orenes2023using]. This issue is addressed in AssertLLM through the application of RAG techniques, which enrich the model with specific knowledge on SVA and Formal Property Verification (FPV). In the evaluation results, GPT-4 failed to generate any correct SVAs for IO ports and only succeeded in creating accurate reset check assertions for registers, leading to an overall correct proportion of only 8%.
@@ V-A Coverage in SVA Evaluation
Some works [orenes2023using] propose to leverage the coverage metric, especially the cone-of-influence (COI) coverage to evaluate the quality of generated SVAs. COI coverage relies on analyzing the signals exercised during simulation or formal verification, which significantly involves internal signals within the design.